a) ∫𝝅 𝒙𝒔𝒊𝒏𝟐𝒙 𝒅𝒙

𝟎

b)∫𝒙𝟐 𝒅𝒙 (𝒙+𝟏)

c)∫ 𝒙+𝟏 𝒅𝒙 𝒙𝟐 −𝟑𝒙+𝟐


1
Expert's answer
2021-05-07T09:01:47-0400

a)


0πxsin(2x)dx\displaystyle\int_{0}^\pi x\sin(2x)dx

xsin(2x)dx\int x\sin(2x)dx

udv=uvvdu\int udv=uv-\int v du

u=x,du=dxu=x, du=dx


dv=sin(2x)dx,v=sin(2x)dx=12cos(2x)dv=\sin(2x)dx, v=\int\sin(2x)dx=-\dfrac{1}{2}\cos(2x)

xsin(2x)dx=12xcos(2x)+12cos(2x)dx\int x\sin(2x)dx=-\dfrac{1}{2}x\cos(2x)+\dfrac{1}{2}\int\cos(2x)dx

=12xcos(2x)+14sin(2x)+C=-\dfrac{1}{2}x\cos(2x)+\dfrac{1}{4}\sin(2x)+C

0πxsin(2x)dx=[12xcos(2x)+14sin(2x)]π0\displaystyle\int_{0}^\pi x\sin(2x)dx=\big[-\dfrac{1}{2}x\cos(2x)+\dfrac{1}{4}\sin(2x)\big]\begin{matrix} \pi \\ 0 \end{matrix}

=12π+0(0+0)=π2=-\dfrac{1}{2}\pi+0-(-0+0)=-\dfrac{\pi}{2}


b)

x2x+1dx=x2+xx1+1x+1dx\int\dfrac{x^2}{x+1}dx=\int\dfrac{x^2+x-x-1+1}{x+1}dx

=xdxdx+1x+1dx=\int xdx-\int dx+\int \dfrac{1}{x+1}dx

=x22x+lnx+1+C=\dfrac{x^2}{2}-x+\ln|x+1|+C

c)

x+1x23x+2dx=x+1(x1)(x2)dx\int\dfrac{x+1}{x^2-3x+2}dx=\int\dfrac{x+1}{(x-1)(x-2)}dx

x+1(x1)(x2)=Ax1+Bx2\dfrac{x+1}{(x-1)(x-2)}=\dfrac{A}{x-1}+\dfrac{B}{x-2}

=A(x2)+B(x1)(x+1)(x2)=\dfrac{A(x-2)+B(x-1)}{(x+1)(x-2)}

x=1:A=2x=1: A=-2

x=2:B=3x=2: B=3

x+1x23x+2dx=21x1dx+31x2dx\int\dfrac{x+1}{x^2-3x+2}dx=-2\int \dfrac{1}{x-1}dx+3\int \dfrac{1}{x-2}dx

=2lnx1+3lnx2+C=-2\ln|x-1|+3\ln|x-2|+C